Lulabalula:

Is the expression 2f + 4f + 2 – 3 equivalent to 6f – 1? 6f – 1 evaluated at f = 9 is 53. 2f + 4f + 2 – 3 evaluated at f = 9 is also 53. 6f – 1 evaluated at f = 3 is 17. What is 2f + 4f + 2 – 3 evaluated at f = 3?

Lulabalula:

@Elsa213

Lulabalula:

help plz

ZombieLover:

We have that the expression 2f + 4f + 2 – 3> we can group it (2f+4f)+(2-3) (2f+4f)+(2-3)=(6f-1) therefore the expression [2f + 4f + 2 – 3] is equivalent to [6f-1] if the expression [6f-1] for f=3 is 17 then the expression [2f + 4f + 2 – 3] for f=3 is also 17 let's check it [2*3 + 4*3 + 2 – 3]> [6+12+2-3]=[20-3]=17> is ok

KjSaif:

@lulabalula

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KjSaif @lulabalula \(\color{#0cbb34}{\text{End of Quote}}\) Hm?

KjSaif:

I was wondering if you saw what @zombielover posted for you

Lulabalula:

yeah sorry